# Fact-Checking with Contextual Narratives: Leveraging Retrieval-Augmented LLMs for Social Media Analysis

### v1

@@ 5.3. Ablation Studies on Clustering
The clustering performance metrics, in Table 4, reveal key insights into the effectiveness of different methods across the real-world datasets 5PilsOOC, Verite, and DP. The evaluation metrics include the average silhouette score (Sil)(Rousseeuw, 1987) (indicating cluster cohesion and separation, with higher values being better) and the Davies-Bouldin(DB)(Davies and Bouldin, 1979) index (where lower values indicate better-defined clusters). Overall, K-means clustering consistently outperforms spectral clustering, with K-means using 8 clusters achieving the highest silhouette scores and lowest Davies-Bouldin indices across all datasets, indicating well-defined and compact clusters. In contrast, spectral clustering struggles, particularly as the number of clusters increases, showing lower silhouette scores and higher DB indices, which suggests less optimal separation. Dataset-specific trends also emerge, with 5PilsOOC exhibiting the best clustering structure, while Verite appears to be the most challenging, as indicated by its lower silhouette scores, especially for spectral clustering. The DP dataset follows a similar trend to 5PilsOOC, reinforcing the overall superiority of K-means. These results suggest that K-means with a higher number of clusters (e.g., 8) is the most effective method for these datasets, while spectral clustering may not be as well suited due to its lower cohesion and separation performance.
5.4. Baseline performance using MLLMs
Table 5 compares Multimodal Large Language Models (MLLMs) with our CRAVE framework on the binary veracity prediction task. For both InstructBLIP (Dai et al., 2023) and LLAVA (Liu et al., 2024a), we use Vicuna-13B (Zheng et al., 2023) as the underlying language model. All models are evaluated on the same input: the claim and the retrieved evidence. However, performance varies significantly due to how they process and reason over noisy, multimodal contexts.
Additionally, unlike MLLMs—whose performance may be skewed by biases from pretraining on large-scale web data—CRAVE is less influenced by domain priors or cultural biases. This is particularly advantageous in datasets like 5PilsOOC and DP, where MLLMs struggle with claims from underrepresented regions or topics. CRAVE’s zero-shot approach and reliance on structured narratives lead to more robust decision-making, regardless of cultural or regional context.
Method Verite MMFakebench 5PilsOOC NewsCLIPpings DP
Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
InstructBLIP(C) 0.50 0.51 0.50 0.52 0.55 0.54 0.55 0.55 0.54 0.54
InstructBLIP(V) 0.63 0.56 0.62 0.59 0.50 0.46 0.50 0.45 0.40 0.34
InstructBLIP(T) 0.56 0.52 0.50 0.51 0.50 0.49 0.47 0.44 0.43 0.40
LLAVA(C) 0.66 0.56 0.68 0.65 0.53 0.38 0.54 0.45 0.40 0.27
LLAVA(T) 0.68 0.63 0.71 0.70 0.60 0.54 0.69 0.67 0.49 0.45
CRAVE 0.82 0.82 0.78 0.76 0.81 0.81 0.85 0.85 0.73 0.73
Table 5. Comparison against MLLMs across datasets. C denotes models using only the Claim. V denotes the version using both images and text as evidence, and T denotes versions using only the text component of the evidence.
6. User Study: Explanation quality and preference
@@
6.1. Study 1: Clarity and comprehensiveness evaluation
To assess the quality of fact-checking explanations produced by our system, we conducted a human evaluation using Amazon Mechanical Turk (AMT). Participants were shown a news story along with a set of retrieved narrative facts, the system-generated veracity prediction, and the system-generated explanation. Their task was to rate the explanation on two dimensions: clarity and comprehensiveness. Clarity refers to how easy the explanation is to understand, and comprehensiveness measures the extent to which the explanation incorporates key facts to justify the verdict. Ratings were provided on a 5-point Likert scale (using discrete values 1, 3, and 5). AMT workers were filtered to include only those who spent more than one minute on the task and had a historical approval rate above 95%; out of 56 participants, 39 met these criteria. Table 6 summarizes the evaluation statistics for both LLM-generated scores and the filtered AMT evaluations. While the participants were not aware of the ground truth, we partitioned the scores for correct and incorrect predictions. The data indicate that explanations for correct predictions generally receive higher clarity scores than those under incorrect predictions. For example, LLM evaluations show an average clarity score of 3.28 for correct predictions versus 2.99 for incorrect ones. AMT evaluations yield higher scores overall. Underscoring the value of human judgment in assessing explanation quality to detect nuances in explanation quality more reliably.
@@ 6.1. Study 1: Clarity and comprehensiveness evaluation
Table 6. Evaluation statistics for CRAVE explanation quality. For each participant (LLM, AMT) and partition (Overall, Correct predictions, Incorrect predictions), the table reports counts for ratings along with the average score for both comprehensiveness and clarity.
6.2. Study 2: Preference comparison of explanations
In the second study, we examine the preference between two types of explanations for a news story: the user-generated explanation (community notes) and the system-generated explanation. The task is to select the explanation that is more intuitive, clear, and concise, and that better supports the story’s ground truth veracity. Table 7 presents the preference fractions from both LLM and AMT workers. While LLM assessments show a preference of 67.11% for the CRAVE-generated explanation compared to 32.89% for the user explanation, AMT workers—based on a sample of 10 workers—exhibited a preference of 57.66% for the fact-checker explanation versus 42.34% for the user-generated notes. These findings suggest that, across both evaluation modalities, the system-generated explanation is consistently viewed as more effective in conveying the veracity of the story.
Source User Explanation (Notes) CRAVE Explanation Workers
@@ 6.2. Study 2: Preference comparison of explanations
Table 7. Preference for explanation amongst LLM and AMT worker: Both LLM and AMT workers prefer CRAVE explanations over Community Notes.
7. Limitations
Figure 4 presents qualitative examples highlighting key factors behind misclassifications, linked to three main limitations of the current framework. First, the model struggles when an image is widely reused in fabricated contexts across multiple sources, allowing the dominant fake narrative to skew retrieval results (e.g., the leftmost mountain landscape). Second, the system treats all sources equally, without factoring in domain credibility—amplifying the influence of low-credibility outlets. Third, while 5W1H-based contextualization helps flag surface-level inconsistencies (e.g., named entities), it misses subtler linguistic cues necessary for disambiguating misinformation, as shown in the second sample. The third and fourth examples illustrate failures in verifying true claims, mainly due to the lack of retrievable evidence—often caused by scraping limitations. In such cases, images are retrieved, but the accompanying text from source sites is inaccessible. Lastly, some errors stem from questionable dataset labels, as seen in the final example.
@@ 7. Limitations
Figure 4. Misclassified samples are color-coded—orange for those labeled ‘True’, blue for those labeled ‘Fake’. A sample with a questionable ground-truth label (‘True’) is shown in grey. Zoom in for details
8. Conclusion
@@
We introduced CRAVE, a novel framework designed to address multimodal misinformation on social media. CRAVE first organizes retrieved evidence into coherent narrative clusters and then employs aa LLM as a judge to assess the veracity of claims based on these narratives. By reasoning over contextualized clusters, CRAVE achieves state-of-the-art performance in zero-shot veracity prediction across multiple datasets, including Verite, MMFakeBench, and 5PilsOOC. Our user study further demonstrates that CRAVE’s explanations are perceived as coherent and comprehensive in many cases, even preferred over community-generated notes. Future directions for CRAVE including improving robustness against pervasively reused images and mitigating occasional issues with subtle linguistic cues.

### v2

@@ V-C4 Study 4 - Cluster method
Overall, K-means clustering consistently outperforms spectral clustering, achieving the higher silhouette scores and lowest Davies-Bouldin indices across all datasets, indicating well-defined and compact clusters. In contrast, spectral clustering struggles suggesting less optimal separation. Dataset-specific trends also emerge, with 5PilsOOC exhibiting the best clustering structure, while Verite appears to be the most challenging, as indicated by its lower silhouette scores, especially for spectral clustering. Interestingly, GPT 4o-mini forms the most coherent clusters, but this is often at the cost of completely ignoring outliers which leads to the lower overall accuracy.
V-D User study: Explanation quality and preference
@@
V-D1 Study 1 - Clarity and comprehensiveness evaluation
To assess the quality of fact-checking explanations produced by our system, we conducted a human evaluation using Amazon Mechanical Turk (AMT). Participants were shown a news story along with a set of retrieved narrative facts, the system-generated veracity prediction, and the system-generated explanation. Their task was to rate the explanation on two dimensions: clarity and comprehensiveness. Clarity refers to how easy the explanation is to understand, and comprehensiveness measures the extent to which the explanation incorporates key facts to justify the verdict. Ratings were provided on a 5-point Likert scale (using discrete values 1, 3, and 5). AMT workers were filtered to include only those who spent more than one minute on the task and had a historical approval rate above 95%; out of 56 participants, 39 met these criteria. Table VII summarizes the evaluation statistics for both LLM-generated scores and the filtered AMT evaluations. While the participants were not aware of the ground truth, we partitioned the scores for correct and incorrect predictions. The data indicate that explanations for correct predictions generally receive higher clarity scores than those under incorrect predictions. For example, LLM evaluations show an average clarity score of 3.28 for correct predictions versus 2.99 for incorrect ones. AMT evaluations yield higher scores overall, underscoring the value of human judgment in assessing explanation quality to detect nuances in explanation quality more reliably.
@@ V-D1 Study 1 - Clarity and comprehensiveness evaluation
TABLE VII: Evaluation statistics for CRAVE explanation quality. For each participant (LLM, AMT) and partition (Overall, Correct predictions, Incorrect predictions), the table reports counts for ratings along with the average score for both comprehensiveness and clarity.
V-D2 Study 2 - Preference comparison of explanations
In the second study, we examine the preference between two types of explanations for a news story: the user-generated explanation (community notes) and the system-generated explanation. The task is to select the explanation that is more intuitive, clear, and concise, and that better supports the story’s ground truth veracity. Table VIII presents the preference fractions from both LLM and AMT workers. While LLM assessments show a preference of 67.11% for the CRAVE-generated explanation compared to 32.89% for the user explanation, AMT workers—based on a sample of 10 workers—exhibited a preference of 57.66% for the CRAVE explanation versus 42.34% for the user-generated notes. These findings suggest that, across both evaluation modalities, the system-generated explanation is consistently viewed as more effective in conveying the veracity of the story.
Source Explanation Workers
Notes (User) CRAVE
@@ V-D2 Study 2 - Preference comparison of explanations
TABLE VIII: Preference for explanation amongst LLM and AMT worker: Both LLM and AMT workers prefer CRAVE explanations over Community Notes.
@@
Figure 5: Misclassified samples are color-coded—orange for those labeled ‘True’, blue for those labeled ‘Fake’. A sample with a questionable ground-truth label (‘True’) is shown in grey. Zoom in for details.
V-E Critical analysis
V-E1 Analysis 1 - Primary causes of errors
Figure 5 presents qualitative examples illustrating key factors behind misclassifications, corresponding to three main types of errors identified in our framework. First, clustering issues occur when an image is widely reused across multiple fabricated contexts, causing the dominant fake narrative to skew evidence grouping and retrieval results (e.g., the leftmost mountain landscape in Figure 5). Second, reasoning errors arise because the system treats all sources equally without weighing domain credibility, amplifying the influence of low-credibility outlets. While 5W1H-based contextualization helps detect surface-level inconsistencies such as named entities, it often misses subtler linguistic cues needed to disambiguate misinformation, as shown in the second sample of Figure 5. Third, retrieval issues contribute to failures in verifying true claims, frequently due to scraping limitations that prevent access to relevant textual evidence even when images are successfully retrieved; this is illustrated by the third and fourth examples in Figure 5. Lastly, some errors stem from questionable dataset labels, as seen in the final example.
Our error analysis reveals three main failure categories, with quantitative results shown in Table IX.
  1. 1.
    Retrieval issues (most common): Errors stem from incomplete retrieval (incom), finding only false evidence (False), or a complete lack of relevant evidence (NO).
  2. 2.
    Clustering issues: True evidence is sometimes grouped with highly similar false narratives (Under-clustering: K$<$). For claims with multiple sub-parts (like in DP dataset), the evidence gets fragmented across clusters (Over-clustering: K$>$).
  3. 3.
    Reasoning issues: The LLM struggles with nuanced actions (Act), missing or vague temporal terms like “recently” (Tmp), and non-Western entities (Ent). A high volume of fake narratives (CN) can also overwhelm its final judgment.
Dataset Label Clustering Reasoning Retrieval
K$<$ K$>$ GT Act CN Ent Tmp False incom NO
5PILS_OOC True 4.4 0.0 9.2 4.8 8.3 5.7 8.7 1.7 20.1 21.4
False 1.3 0.0 0.9 2.2 3.9 0.0 1.7 5.2 0.4 0.0
DP True 12.1 3.8 5.7 13.4 12.1 7.6 3.8 0.0 24.2 12.1
False 0.0 0.0 0.6 0.0 1.3 0.0 0.0 3.2 0.0 0.0
TABLE IX: Error Type Distribution by Dataset and Label (%)
V-E2 Analysis 2 - Evidence retrieval failure
The “not enough data” (NoD) predictions occur when our retrieval pipeline fails to find sufficient evidence—often due to niche topics, deepfakes, or paywalled content. In our binary classification task, we conservatively label NoD as “Misleading.” These predictions are thus adjudged incorrect when the Ground Truth (GT) is True, and correct when GT is False. In Table X we present our findings. High NoD rates in NewsCLIPpings result from using only the dataset-provided evidence, with no retrieval refinement. While our method does not directly detect deepfake images through image analysis, we address such cases indirectly—either when the content is flagged as deepfake by external sources or when it lacks presence in reputable news outlets. Consequently, in the MMFakeBench dataset, fake content involving deepfake images—which cannot be corroborated through external evidence retrieval—results in the highest NoD percentage ($9.71$).
Dataset % Total Pred False Pred which NoD Pred with NoD Pred with
Name as False are NoD GT True GT False
DP 59.8 5.5 2.75 2.75
NewsCLIPpings 61.9 10.3 4.22 6.11
MMFakeBench 81.7 14.8 5.06 9.71
5PILS_OOC 63.1 3.3 1.80 1.47
Verite 69.9 2.1 0.99 1.11
TABLE X: Analysis of “Not Enough Data” Model Predictions Across Datasets
VI Conclusion
@@
We introduced CRAVE, a novel framework designed to address multimodal misinformation on social media. CRAVE first organizes retrieved evidence into coherent narrative clusters and then employs an LLM as a judge to assess the veracity of claims based on these narratives. By reasoning over contextualized clusters, CRAVE achieves state-of-the-art performance in zero-shot veracity prediction across multiple datasets, including Verite, MMFakeBench, and 5PilsOOC. Our user study further demonstrates that CRAVE’s explanations are perceived as coherent and comprehensive in many cases, even preferred over community-generated notes.
VII Limitations and Future Work
While CRAVE demonstrates a significant step forward in multimodal fact-checking, we acknowledge several limitations that define its current scope and provide clear directions for future research.
  1. 1.
    Dependencies and Source Trust: The framework’s performance is fundamentally bound by the capabilities and biases of its external dependencies: search engines and Large Language Models. Our reliance on commercial search APIs means the retrieved evidence pool is subject to uncontrollable algorithmic biases (e.g., geographical or political), and we currently treat all retrieved sources with equal weight, making the system susceptible to well-optimized, low-credibility content. This underscores the need for future work in source credibility modeling to be integrated into the evidence-gathering phase.
  2. 2.
    Pragmatic Heuristics: Our methodology incorporates pragmatic heuristics to ensure broad applicability, most notably the choice of a static K=4 for clustering. Our own analysis (Figure 4) powerfully illustrates that the optimal cluster count is claim-dependent. We therefore present K=4 not as a universal optimum, but as a necessary and empirically grounded baseline for a dataset-agnostic system.
  3. 3.
    Computational Cost and Applicability: The comprehensive nature of our multi-step pipeline introduces a significant computational cost and latency ($\sim$4 minutes serially, $\sim$1 minute in batches). This positions CRAVE not as a tool for real-time content moderation, but rather for deep, offline analyses where rigor is prioritized over speed, such as in journalistic investigations, academic research, or platform policy audits.
  4. 4.
    Scope of Modalities: Finally, our current work is scoped to text-and-image verification. Future iterations should aim to expand the framework’s capabilities to address other critical modalities of misinformation, such as video and audio, to provide a more holistic solution to the information forensics challenge.
VIII Ethics statement
While our explanation explicitly cites the narratives used with the narrative themselves being annotated with sources used, we emphasize that CRAVE is a decision-support, not an autonomous arbiter of truth, and must be used with an awareness of its potential for error.